### v1

# Agentic-R1: Distilled Dual-Strategy Reasoning

Abstract
Current long chain-of-thought (long-CoT) models excel at mathematical reasoning but rely on slow and error-prone natural language traces. Tool-augmented agents address arithmetic via code execution, but often falter on complex logical tasks. We introduce a fine-tuning framework, DualDistill, that distills complementary reasoning strategies from multiple teachers into a unified student model. Using this approach, we train Agentic-R1, which dynamically selects the optimal strategy for each query, invoking tools for arithmetic and algorithmic problems, and using text-based reasoning for abstract ones. Our method improves accuracy across a range of tasks, including both computation-intensive and standard benchmarks, demonstrating the effectiveness of multi-strategy distillation in achieving robust and efficient reasoning. Our project is available at https://github.com/StigLidu/DualDistill.
Agentic-R1: Distilled Dual-Strategy Reasoning
Weihua Du Pranjal Aggarwal Sean Welleck Yiming Yang Language Technologies Institute, Carnegie Mellon University {weihuad, pranjala, swelleck, yiming}@cs.cmu.edu
1 Introduction
A recently proposed reasoning paradigm for language models, long chain-of-thought (long-CoT) reasoning, has achieved state-of-the-art performance on challenging tasks such as mathematical problem solving Guo et al. (2025); Jaech et al. (2024). By allocating a large inference budget, these models generate reasoning trajectories with iterative self-verification and refinement. Despite this progress, open-source long-CoT models remain limited: Their reasoning traces rely solely on natural language, which is both computationally expensive and error-prone without explicit verification.
In contrast, tool-aided reasoning provides greater efficiency and reliability, particularly for large-scale numerical computations and tasks that require rigorous verification Gao et al. (2023). Advanced agent frameworks, such as OpenHands Wang et al. (2024), place language models in a multi-turn environment with a code interpreter and other tools. The resulting agentic trajectories are effective for tool-intensive tasks, but often fall short on abstract or conceptually complex reasoning problems Duan et al. (2024).
To leverage the strengths of both reasoning and tool-based strategies, we introduce DualDistill, a novel distillation framework (Fig. 1) that combines trajectories from two complementary teachers: one reasoning-oriented, the other tool-augmented, in a unified student. The resulting model, Agentic-R1, learns to mix both strategies and dynamically selects the most appropriate one for each problem, executing code for arithmetic and algorithmic tasks and reasoning in natural language for abstract ones. Furthermore, the student can continue to improve via self-distillation, better calibrating its strategy selection to its actual capabilities. Our contributions are as follows.
DualDistill, a distillation framework that enables a language model to learn from multiple teacher models with complementary capabilities through solution trajectory composition.
Agentic-R1, a distilled student model that achieves strong performance in mathematical tasks requiring both tool use and complex reasoning.
2 Related Work
While prior efforts have integrated external tools into language models Gao et al. (2023); Schick et al. (2023); Nakano et al. (2022), they are often specialized to either non-math domains or are confined to shorter reasoning chains. Concurrently, the paradigm of long chain-of-thought (long-CoT) reasoning or inference-time compute has demonstrated significant improvements Guo et al. (2025); Jaech et al. (2024). However, these approaches can be difficult to control and may suffer from ‘overthinking’, particularly when applied to tool-use scenarios Cuadron et al. (2025). Recent works have combined tool use with long CoT reasoning Feng et al. (2025); Song et al. (2025), but these are often applied to different domains or rely on reinforcement learning, which can be less stable than our proposed distillation method. To the best of our knowledge, DualDistill is the first framework to employ distillation with trajectory composition from two heterogeneous teacher models, one specializing in agentic tool-use and the other in pure textual reasoning, creating a unified student model capable of adaptively leveraging both strategies. See Appendix B for a more detailed discussion.
3 Method
As illustrated in Fig. 1, DualDistill uses trajectory composition to distill the knowledge of the complementary teachers to the student model. The student model then applies self-distillation for a deeper understanding of the strategies.
3.1 Trajectory Composition
Let be a training set, where denotes the -th problem and is its corresponding reference answer. Let and be two distinct teacher policies, where represents the agentic teacher and the reasoning teacher. For each training instance , we randomly select the initial teacher by sampling a binary indicator and then produce solutions and as follows:
That is, after one teacher generates the initial solution , the other teacher subsequently generates the second solution , conditioned on both the original problem and the preceding solution .
We evaluate the correctness of each solution using a rule-based grader, assigning binary correctness scores to and , respectively. The distilled training trajectories are then composed based on these correctness scores.
: The first teacher produces an incorrect solution, and the second teacher successfully corrects it. The composed trajectory is structured as . (Here denotes concatenation and is a transition segment, described later).
: Both teachers provide correct solutions. We create a trajectory to reflect complementary correct strategies.
: Only the initial teacher provides a correct solution. In this scenario, the composed trajectory includes only .
: Both teachers do not solve the problem correctly. In this case, we just discard the problem without composing any trajectory.
The transition segments and are predefined sentences indicating strategy shifts (e.g., "Wait, using text reasoning is too tedious, let us try code reasoning."). More examples and detailed descriptions can be found in Appendix A.4.1.
3.2 Training Instance Selection
We curate a training set with the instances for which one strategy has a clear advantage over the other in performance. Using an existing dataset such as GSM8K Cobbe et al. (2021) would be insufficient in this sense as most of the problems are relatively simple and can be solved by either strategy without a significant performance difference. Instead, we construct two contrasting subsets of Math problems from DeepMath He et al. (2025): one can benefit more from tool-assisted reasoning, while the other can benefit more from pure text-based reasoning. After composition, we apply additional filtering to balance the training dataset, resulting in k distilled trajectories. Detailed statistics can be found in Appendix A.3.2. Further filtering details are provided in the Appendix A.3.1.
3.3 Teacher and Student Models
As the teacher of agentic reasoning, we utilize OpenHands Wang et al. (2024), a tool-assisted agent built upon Claude-3.5-Sonnet Anthropic (2024) to employ human-designed problem-solving pipelines. As the teacher of text-based reasoning, we adopt Deepseek-R1 Guo et al. (2025). The details can be found in Appendix A.4.2.
As for the student model, we adopt Deepseek-R1-Distill-7B, which has been fine-tuned on pure text-based reasoning trajectories and also exposed to code-related data during pretraining. We deliberately choose a model already familiar with both modalities to minimize the amount of training data required for the strategic composition. We want to examine whether it can effectively learn multiple problem-solving strategies.
3.4 Self-Distillation
Although the student model learns problem-solving strategies from multiple teachers, it can still underperform compared to them due to limitations such as the smaller model sizes, leading to an ability mismatch. For instance, we find that the student sometimes uses tools for problems that could be solved more reliably through simple reasoning. While this approach is valid, it can introduce errors because the student’s tool-use ability is less mature than that of the teachers, occasionally leading to incorrect code and wrong answers.
To address it, we introduce self-distillation to help the student further refine its strategy selection based on its capabilities and the given problem. Our self-distillation process involves the student model generating candidate solutions, with teacher models providing verification or corrections as auxiliary supervision. The process reinforces effective strategies and corrects suboptimal ones. Specifically, given a training set and the student policy finetuned on distillation data from teachers , we sample trajectories for each problem :
We then apply a binary grader to evaluate trajectory accuracy. Let be the score of the -th trajectory and be the average score for , i.e.,
If , the student cannot fully solve problem , and we collect informative trajectories from its output to form a self-distillation buffer for further training. Specifically:
If , we add a correct trajectory generated by the student, followed by a verification from a teacher model, to the replay buffer;
If , we add an incorrect student trajectory, along with a corrected solution provided by a teacher, to the buffer.
Here, and are hyperparameters that control the difficulty of the problems selected for the replay buffer. We set and , a relatively low threshold that encourages diversity in the collected examples. In addition, we use trajectory samples per problem. Verification (or correction) consists of a correct trajectory from the teacher model with some transition words; see Appendix A.4.1 for details. Because we observed a gap in the coding ability between the student and the teacher, we provide only text-based reasoning solutions as the teacher’s answers at this stage.
The pseudocode for our complete algorithm, DualDistill, is listed in Algorithm 1.
4 Experiments
| Model | Budget | DeepMath-L | Combinatorics300 | MATH500 | AIME | AMC | avg. |
| Qwen2.5-7B-Instruct (w/o tool) | S L | 17.2 17.5 | 21.8 21.8 | 75.1 75.2 | 8.0 8.0 | 42.9 42.9 | 33.0 33.1 |
| Qwen2.5-7B-Instruct (w/ tool) | S L | 34.7 34.7 | 28.9 28.9 | 70.2 70.2 | 14.7 14.7 | 51.1 51.1 | 39.9 39.9 |
| DeepSeek-R1-Distill-7B | S L | 34.7 56.3 | 34.7 44.5 | 83.1 89.2 | 23.3 40.7 | 61.2 84.8 | 47.4 63.1 |
| Agentic-R1-7B (ours) | S L | 37.0 59.3 | 36.9 49.4 | 80.0 82.4 | 28.0 40.7 | 64.3 82.2 | 49.3 62.8 |
| Agentic-R1-7B-SD (ours) | S L | 40.0 65.3 | 38.2 52.0 | 82.5 93.3 | 27.3 40.7 | 66.3 85.8 | 50.9 67.4 |
4.1 Benchmarks
We evaluated our method on several benchmarks that test different aspects of mathematical reasoning, including tasks where tool-aided calculation is hypothesized to provide a significant advantage.
DeepMath-L.
DeepMath He et al. (2025) is a comprehensive dataset of mathematical and STEM problems compiled from various benchmarks. We curate a subset of 87 problems with large answers (absolute value greater than ). These problems are excluded from our fine-tuning data, although they may appear in some pretraining corpora. We refer to this evaluation set as DeepMath-L, with the assumption that code-aided computation is more effective in solving such problems.
Combinatorics300.
This benchmark consists of 300 combinatorics problems aggregated from diverse math test sets. Each problem yields an answer larger than , reflecting the factorial growth in combinatorial counts. We hypothesize that tool-aided reasoning is important for handling the enumeration and sampling required in such tasks.
Standard Mathematical Benchmarks.
4.2 Baselines
We compare against the following strong baselines:
DeepSeek-R1-Distill. A distilled version of DeepSeek-R1 fine-tuned on long chain-of-thought trajectories, representing a strong baseline for pure language-based reasoning.
Qwen-2.5-Instruct (w/ tool, w/o tool) Yang et al. (2024). A general-purpose short-CoT model with optional tool-use capabilities. The tool-augmented variant serves as a competitive baseline for tool-aided strategies.
The training configuration details are provided in Appendix A.2.
4.3 Evaluation Metrics
To assess both reasoning quality and computational efficiency, we adopt the Accuracy at Budget metric. Let be the trajectory generated by the model, where each denotes the -th output token, and let be the reference answer. The accuracy under a computational budget is defined as:
where G is a binary grader that evaluates whether the model’s partial output matches the ground truth. We report results under two budgets: Standard (S, 4096), a moderate token budget for language model reasoning, and Large (L, 32768), which approximates an unbounded budget and allows the model to reason adequately. Inference and grader details can be found in Appendix A.6.
4.4 Results
As shown in Table 1, our student model, Agentic-R1, demonstrates substantial performance improvements in DeepMath-L and Combinatorics300, two challenging datasets that benefit from both agentic and reasoning strategies. Specifically, our model outperforms two similarly sized models, each specializing exclusively in tool-assisted (Qwen2.5-7B-Instruct) or pure reasoning (Deepseek-R1-Distill-7B) strategies. Agentic-R1 surpasses tool-based models by intelligently adopting reasoning strategies when appropriate, while maintaining greater efficiency compared to pure reasoning models on standard mathematical tasks. However, we note a slight performance decrease in relatively simpler benchmarks (MATH500) compared to the pure text-reasoning model, and a detailed discussion is provided in the limitations section.
Qualitative Examples.
We provide illustrative trajectories demonstrating Agentic-R1’s adaptive strategy-switching capability: (1) initially using the tool-assisted strategy and then switching to textual reasoning to correct an incorrect initial solution (Fig. 6); and (2) starting with textual reasoning and then switching to the tool-assisted strategy to bypass tedious manual calculations (Fig. 7).
Agentic-R1 Knows When to Use Tools.
An intriguing observation is that Agentic-R1 learns when to appropriately invoke code tools purely through supervised fine-tuning. For instance, Combinatorics300 contains problems involving large numerical computations, which makes the tools particularly beneficial. Consequently, Agentic-R1 activates at least one code execution tool in of the Combinatorics300 problems, while the usage of the tool drops to only in the relatively simpler AMC dataset.
Agentic-R1 Learns from Imperfect Teachers.
Although OpenHands, based on Claude-3.5-Sonnet, is not a strong standalone reasoning agent and sometimes performs worse than the student’s initial model (R1-Distill-7B), the student model still effectively acquires valuable agentic strategies through distillation. For example, the agentic strategy teacher achieves only accuracy on Combinatorics300, yet after training, the student’s performance improves significantly from to , surpassing the teacher. This shows that demonstrations from an imperfect agentic teacher can still yield meaningful gains in the student.
4.5 Ablation Study
| Dataset | DeepMath-L | AIME | AMC |
| w/o composition | 40.0% | 34.0% | 50.8% |
| w/ composition | 59.3% | 40.7% | 82.2% |
Trajectory Composition.
To verify the effectiveness of our data composition strategy, we compare it with a training strategy that does not use composition, meaning that each student trajectory is either fully generated by the agentic teacher or fully generated by the reasoning teacher. As shown in Table 2, our composition strategy consistently surpasses its non-composition counterpart.
5 Conclusion
We propose DualDistill, an efficient distillation framework based on trajectory composition, allowing a student model to learn from multiple teacher models specialized in different domains of problem solving. Using the appropriate strategy for each problem, our trained model, Agentic-R1, achieves superior performance in benchmarks that require both reasoning and tool-assisted capabilities. This approach demonstrates the potential for unifying diverse problem-solving strategies within a single model, opening new directions for building versatile and adaptive language agents.
Limitations
While our approach demonstrates strong overall performance, several limitations remain for future work. First, the transition words connecting different strategies within composed trajectories are currently designed manually. As a result, the output produced by the trained student model can occasionally lack naturalness and fluidity, especially when switching between strategies. Developing methods for more coherent and automatic transitions of strategies could further enhance the readability and overall quality of the content generated by the student model.
Second, our training dataset contains approximately k composed trajectories. While this appears sufficient to teach a model that has been pre-trained on both text reasoning and code generation (e.g., Deepseek-R1-Distill-7B) to choose between strategies, it is likely insufficient for training a model to learn a new reasoning strategy from scratch. For example, Deepseek-R1-Distill was fine-tuned on over k distilled examples to acquire long CoT reasoning capabilities. Expanding the dataset and covering a wider range of strategies will be an important direction for future research.
Appendix A Appendix
A.1 Code and Dataset
Our code is available at https://github.com/StigLidu/DualDistill; Training data is available at https://huggingface.co/datasets/VanishD/DualDistill.
A.2 Training Configuration
Loss Masking.
To prevent the student model from learning incorrect patterns from unsuccessful attempts, we exclude specific segments of trajectories from the loss calculation. Specifically, trajectory segments occurring before a transition from incorrect to correct reasoning (i.e., ) are omitted. Additionally, the executor’s feedback and the code blocks resulting in execution errors are also excluded from influencing the loss computation.
Hyperparameters.
For finetuning the student model Agentic-R1 on teacher distilled trajectories, we use 4 A6000 GPUs for a total of 12.7 hours. The model is trained for 4 epochs using the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of . We set the maximum context length to 16,384 tokens for teacher distillation and 8,192 for self-distillation, and discard all training examples that exceed this limit.
A.3 Dataset Details
A.3.1 Problem Filtering Heuristics
To curate a training dataset that can guide a student model in learning when to apply agentic versus pure text-based reasoning, we construct two subsets of mathematical problems.
Agentic-Favored Subset.
We identify problems where tool use is highly beneficial using two heuristics:
Numerical Scale: Problems whose final integer answers exceed an absolute value of often require nontrivial arithmetic operations or algorithms that are more suitable for tool-assisted computation.
Difficulty Under Constraints: We use a baseline text reasoning-only model, Deepseek-R1-Distill-7B, with a limited context length (4096 tokens). Problems unsolvable under the constraint with one trail are deemed more difficult and suitable for agentic strategies.
Pure Reasoning-Favored Subset.
To balance the dataset, we include problems in which agent execution is error-prone. These are selected by identifying the cases where the tool-assisted strategy fails and produces incorrect output.
We apply this selection process to DeepMath-103K He et al. (2025) and balance the two subsets to ensure that the model sees roughly equal representation from both strategies during training.
A.3.2 Dataset Scale
| type | |||
| number | 685 | 600 | 1393 |
After running the two teachers on the filtered subset and compositing the trajectories, the final distilled dataset contains examples. The detailed count for each correctness category is listed in Tab. 3.
A.4 Composition Trajectory
A.4.1 Transition Segment
When the teacher changes, a hand-designed transition segment is added to signify and point out the meaning of the transition. There are three typical transition segments , which are shown in Table 4. The transition segments used in self-distillation are the same as those used in teacher distillation.
| Meaning | Content |
| tool () text () | Wait, the code is not correct, let’s try text reasoning. |
| text () tool () | Wait, use text reasoning is too tedious, let’s try code reasoning. |
| A () B () | Wait, we can also use {B}-reasoning as an alternative way to verify the solution. |
A.4.2 Trajectory Composition Implementation
To transform multiturn agentic trajectories from OpenHands logs into a suitable training format, we extract content from the log fields labeled ‘thought’, ‘code’, and ‘final thought’ along with their associated executor feedback if any. Each extracted content is then enclosed within distinct resource tags—<think></think>, <code></code>, <answer></answer> or <executor></executor>—and concatenated sequentially. For reasoning trajectories from Deepseek-R1, we specifically apply the <answer></answer> tag to the content outside the long CoT part (i.e., beyond the <think></think> segment).
We aim for the student model to select the most efficient strategy inherently. Thus, we enforce a token budget on the first teacher’s inference: If does not complete within a randomly determined inference budget , the inference is stopped and labeled unsuccessful. In contrast, we do not impose any token budget constraint on the trajectory of the second teacher .
During preliminary experiments, we observed substantial differences in the distributional characteristics between the OpenHands trajectories () and Deepseek-R1 trajectories (). To avoid performance degradation of due to potential contamination from combined inputs, we assume conditional independence and explicitly define the teacher model inference policy as .
A.5 Qualitative Example
We observed that Agentic-R1 shows several promising behaviors: (1) The model initially adopts tool-aided reasoning, but results in wrong results after several attempts, and then the model automatically switches to text reasoning and finally derives the correct answer (Fig. 6); (2) The model initially tries to apply text reasoning for a combinatorial problem, and then changes to tool-aided reasoning to reduce computational complexity (Fig. 7).
A.6 Inference Details
For all evaluation experiments, we use the VLLM framework Kwon et al. (2023) to enable fast inference via prefix caching, which significantly accelerates multi-turn tool calling. In the tool-augmented setting, the language model is allowed to invoke a Python executor up to times per problem, with each execution capped at seconds. During inference, whenever the model outputs the special token </code>, the generation process is temporarily paused, the preceding code block is executed, and the resulting feedback is appended to the ongoing generation enclosed with <executor> </executor> before resuming inference. Although tool execution introduces up to seconds of additional runtime per query, this cost is relatively small compared to the time-intensive pure text reasoning process, which can take several minutes to reach a conclusion using Deepseek-R1-Distill-7B on 2A6000 GPUs. Additionally, the prompt template is listed in Fig. 2.
The grader evaluates output trajectories in two steps:
Exact match: The grader extracts the final non-think block: a code result (‘<executor>…</executor>’) or a direct answer (‘<ans>…</ans>’), and compares it with the gold answer.
Fuzzy match: If no exact match is found, the full output is passed to MathVerify HuggingFace (2025), an open-source verifier that checks mathematical equivalence. This helps capture correct answers that may appear earlier in the trace, ensuring a fairer comparison for long-CoT baselines (e.g., DeepSeek-R1-Distill) when facing length truncation.
A.7 Full Results
We report the performance trend of different models tested in various token budgets. Please refer to Fig. 4 for individual benchmarks and Fig. 3 for the average.
Appendix B Related Work
Tool-Augmented Reasoning.
Integrating external tools into the language model chain-of-thoughts has substantially improved the accuracy of numerical and factual tasks. Early program-aided methods such as PaL Gao et al. (2023) and PoT Chen et al. (2023) demonstrated significant gains by converting reasoning steps into executable programs, delegating precise computations to code interpreters. Other lines of work, including WebGPT Nakano et al. (2022) and ReAct Yao et al. (2023), introduced agent-like reasoning frameworks that interleave tool invocation (e.g., web searches or API calls) within multi-step reasoning. Toolformer Schick et al. (2023) further generalized this approach by training language models to self-supervise API calls on various tasks such as arithmetic, translation, and retrieval. However, unlike DualDistill, these methods typically rely on short chain-of-thoughts primarily using prompting or heuristic-based tool invocation, lacking mechanisms for automatically balancing pure long reasoning against tool use based on task complexity.
Long Chain-of-Thought Reasoning.
Recent approaches have highlighted significant performance improvements by scaling inference-time chain-of-thought length. OpenAI’s O1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) used outcome-driven reinforcement learning to generate extensive reasoning trajectories, substantially outperforming shorter-CoT baselines on complex math and reasoning benchmarks. Similarly, S1 Muennighoff et al. (2025) and L1 Aggarwal and Welleck (2025) demonstrated scaling curves showcasing a log-linear relationship between performance and inference compute. Empirical evidence supports that an increase in inference computation can often yield more cost-effective gains than an increase in model size alone Wu et al. (2025). Nonetheless, long-CoT models frequently encounter overthinking, i.e, generating overly long reasoning that leads to redundant or incorrect outcomes, especially in tool use scenarios, in a phenomenon known as the reasoning-action dilemma Cuadron et al. (2025). Our work addresses these issues by teaching a student model when to switch between internal reasoning and tool-based execution adaptively.
Reasoning Models with Tool-Calling.
Recently, some works have explored the idea of combining long-form reasoning with explicit tool invocation. R1-Searcher Song et al. (2025) and Search-R1 Jin et al. (2025) introduced reinforcement-learning-based retrieval policies within reasoning loops, achieving substantial performance improvements on open-domain question-answering tasks. However, unlike these methods, DualDistill is specifically tailored for math tasks. Similarly, ReTool Feng et al. (2025) trained a reasoning model with tool use for math tasks. However, unlike these approaches that rely on expensive and unstable reinforcement learning techniques, DualDistill is a simple distillation approach, leading to a more data-efficient and practical training setup.
Distillation in Large Language Models.
Knowledge distillation is widely used to transfer capabilities from larger models to smaller and more efficient models Sanh et al. (2019); Hsieh et al. (2023). Recent extensions include multi-teacher distillation frameworks, which aggregate knowledge from multiple similarly structured teachers Li et al. (2024). However, existing distillation works typically assume homogeneous teacher models or single-modal reasoning paradigms. In contrast, our proposed DualDistill explicitly utilizes heterogeneous teacher models: One specialized in agentic tool use and the other in pure textual reasoning. DualDistill proposes an innovative approach to composing trajectories that effectively guides the student to learn from and combine both strategies.
Self-Taught Learning.
A growing line of work explores methods for language models to automatically improve from their own generated outputs or feedback, a paradigm often referred to as self-taught learning or self-distillation. Early approaches in this area explored self-taught learning with pseudo-labels, where a model generates rationales or answers on unlabeled data and subsequently fine-tunes itself on these predictions Zelikman et al. (2022); Wang et al. (2022). Shen et al. (2022) introduces self-distillation by using the outputs of the previous mini-batch as soft targets, improving output consistency without external teachers. In the LLM era, fine-tuning LLMs on self-generated trajectories with a rule-based verifier has been applied in many domains, including coding tasks Jiang et al. (2024), formal proving Lin et al. (2024), math tasks Kumar et al. (2024), and others. A concurrent work Zheng et al. (2025) enables self-evolving reasoning across multiple modalities and fuses them during inference.
Appendix C License
Our training dataset is constructed based on existing datasets, language models, and software. The following lists the relevant resources and their corresponding licenses.
Openhands: An open-source agent framework under the MIT License;
Deepseek-R1: An open-source language model under the MIT License;
Claude-3-5-Sonnet: A commercial language model under a proprietary license; accessible via the Anthropic API and supported cloud platforms;
MATH500: An open-source math dataset under the MIT License;
DeepMath: An open-source math dataset under the MIT License.
All third-party resources were used following their licenses and intended use, as publicly specified. Our model and dataset are released under the MIT License.
Appendix D Use of AI Assistants
For this project, we use AI assistants for paper editing (e.g., grammar, spelling, word choice) and to assist with the running of experiments (e.g., scripting, automation support).
Appendix E Potential Risks
Our work focuses on a model distillation framework for mathematical reasoning tasks, using synthetic and publicly available data. It does not involve deployment or the use of personal data. Therefore, we do not identify significant risks related to safety, privacy, fairness, or security within the current scope.
One potential data-related risk arises from the distillation of the teacher models. Although our training data is public, trajectory distillation may produce misleading or hallucinatory content. Such distilled data may require filtering if applied to sensitive domains in future work.

### v2

@@
 2 Related Work
-While prior efforts have integrated external tools into language models Gao et al.
+Although prior efforts have integrated external tools into language models Gao et al.
 (2023); Schick et al.
 (2023); Nakano et al.
-(2022), they are often specialized to either non-math domains or are confined to shorter reasoning chains.
+(2022), they are often specialized to non-math domains or are confined to shorter reasoning chains.
 Concurrently, the paradigm of long chain-of-thought (long-CoT) reasoning or inference-time compute has demonstrated significant improvements Guo et al.
@@
 (2025).
-Recent works have combined tool use with long CoT reasoning Feng et al.
+Recent works have combined tool use with long-CoT reasoning Feng et al.
 (2025); Song et al.
@@
 The process reinforces effective strategies and corrects suboptimal ones.
-Specifically, given a training set and the student policy finetuned on distillation data from teachers , we sample trajectories for each problem :
+Specifically, given a training set and the student policy fine-tuned on distillation data from teachers , we sample trajectories for each problem :
 We then apply a binary grader to evaluate trajectory accuracy. Let be the score of the -th trajectory and be the average score for , i.e.,
@@
 4 Experiments
-| Model | Budget | DeepMath-L | Combinatorics300 | MATH500 | AIME | AMC | avg. |
-| Qwen2.5-7B-Instruct (w/o tool) | S L | 17.2 17.5 | 21.8 21.8 | 75.1 75.2 | 8.0 8.0 | 42.9 42.9 | 33.0 33.1 |
-| Qwen2.5-7B-Instruct (w/ tool) | S L | 34.7 34.7 | 28.9 28.9 | 70.2 70.2 | 14.7 14.7 | 51.1 51.1 | 39.9 39.9 |
-| DeepSeek-R1-Distill-7B | S L | 34.7 56.3 | 34.7 44.5 | 83.1 89.2 | 23.3 40.7 | 61.2 84.8 | 47.4 63.1 |
-| Agentic-R1-7B (ours) | S L | 37.0 59.3 | 36.9 49.4 | 80.0 82.4 | 28.0 40.7 | 64.3 82.2 | 49.3 62.8 |
-| Agentic-R1-7B-SD (ours) | S L | 40.0 65.3 | 38.2 52.0 | 82.5 93.3 | 27.3 40.7 | 66.3 85.8 | 50.9 67.4 |
+| Model | Budget | DeepMath-L | Combinatorics300 | MATH500 | AIME | AMC | avg. | ||||||||||||||
+| Qwen2.5-7B-Instruct (w/o tool) |
+| Qwen2.5-7B-Instruct (w/ tool) |
+| DeepSeek-R1-Distill-7B |
+| Agentic-R1-7B (ours) |
+| Agentic-R1-7B-SD (ours) |
 4.1 Benchmarks
@@
 As a result, the output produced by the trained student model can occasionally lack naturalness and fluidity, especially when switching between strategies.
-Developing methods for more coherent and automatic transitions of strategies could further enhance the readability and overall quality of the content generated by the student model.
+Moreover, the student model after self-distillation may exhibit repetitive behavior.
+Developing methods for more coherent and automatic transitions between strategies could further enhance the readability and overall quality of the content generated by the student model.
 Second, our training dataset contains approximately k composed trajectories.
@@
 Expanding the dataset and covering a wider range of strategies will be an important direction for future research.
+Acknowledgments
+This work was supported in part by the National Science Foundation under Grant Nos. DMS-2434614 and DMS-2502281.
 Appendix A Appendix
@@
 Hyperparameters.
-For finetuning the student model Agentic-R1 on teacher distilled trajectories, we use 4 A6000 GPUs for a total of 12.7 hours.
+For fine-tuning the student model Agentic-R1 on teacher distilled trajectories, we use 4 A6000 GPUs for a total of 12.7 hours.
 The model is trained for 4 epochs using the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of .
@@
 Difficulty Under Constraints: We use a baseline text reasoning-only model, DeepSeek-R1-Distill-7B, with a limited context length (4096 tokens).
-Problems unsolvable under the constraint with one trail are deemed more difficult and suitable for agentic strategies.
+Problems unsolvable under the constraint with one trial are deemed more difficult and suitable for agentic strategies.
 Pure Reasoning-Favored Subset.
@@
 A.3.2 Dataset Scale
-| type | |||
+| case | |||
 | number | 685 | 600 | 1393 |
-After running the two teachers on the filtered subset and compositing the trajectories, the final distilled dataset contains examples.
+After running the two teachers on the filtered subset and composing the trajectories, the final distilled dataset contains examples.
 The detailed count for each correctness category is listed in Tab.
@@
 A.4.2 Trajectory Composition Implementation
-To transform multiturn agentic trajectories from OpenHands logs into a suitable training format, we extract content from the log fields labeled ‘thought’, ‘code’, and ‘final thought’ along with their associated executor feedback if any.
+To transform multi-turn agentic trajectories from OpenHands logs into a suitable training format, we extract content from the log fields labeled ‘thought’, ‘code’, and ‘final thought’ along with their associated executor feedback, if any.
 Each extracted content is then enclosed within distinct resource tags—<think></think>, <code></code>, <answer></answer> or <executor></executor>—and concatenated sequentially.
@@
 A.5 Qualitative Example
-We observed that Agentic-R1 shows several promising behaviors: (1) The model initially adopts tool-aided reasoning, but results in wrong results after several attempts, and then the model automatically switches to text reasoning and finally derives the correct answer (Fig.
+We observed that Agentic-R1 shows several promising behaviors: (1) The model initially adopts tool-aided reasoning, but yields incorrect outputs after several attempts, and then the model automatically switches to text reasoning and finally derives the correct answer (Fig.
 6); (2) The model initially tries to apply text reasoning for a combinatorial problem, and then changes to tool-aided reasoning to reduce computational complexity (Fig.
@@
 For all evaluation experiments, we use the VLLM framework Kwon et al.
-(2023) to enable fast inference via prefix caching, which significantly accelerates multi-turn tool calling.
+(2023) to enable fast inference via prefix caching, which significantly accelerates multi-turn tool calls.
 In the tool-augmented setting, the language model is allowed to invoke a Python executor up to times per problem, with each execution capped at seconds.
 During inference, whenever the model outputs the special token </code>, the generation process is temporarily paused, the preceding code block is executed, and the resulting feedback is appended to the ongoing generation enclosed with <executor> </executor> before resuming inference.
-Although tool execution introduces up to seconds of additional runtime per query, this cost is relatively small compared to the time-intensive pure text reasoning process, which can take several minutes to reach a conclusion using Deepseek-R1-Distill-7B on 2A6000 GPUs.
+Although tool execution introduces up to seconds of additional runtime per query in the worst case, this cost is relatively small compared to the time-intensive pure text reasoning process, which can take several minutes to reach a conclusion using DeepSeek-R1-Distill-7B on 2A6000 GPUs.
 Additionally, the prompt template is listed in Fig.
@@
 The grader evaluates output trajectories in two steps:
-Exact match: The grader extracts the final non-think block: a code result (‘<executor>…</executor>’) or a direct answer (‘<ans>…</ans>’), and compares it with the gold answer.
+Exact match: The grader extracts the final non-think block: a code result (‘<executor>…</executor>’) or a direct answer (‘<answer>…</answer>’), and compares it with the gold answer.
 Fuzzy match: If no exact match is found, the full output is passed to MathVerify HuggingFace (2025), an open-source verifier that checks mathematical equivalence.
@@
 Tool-Augmented Reasoning.
-Integrating external tools into the language model chain-of-thoughts has substantially improved the accuracy of numerical and factual tasks.
-Early program-aided methods such as PaL Gao et al.
+Integrating external tools into the language model chain-of-thought (CoT) has substantially improved the accuracy of numerical and factual tasks.
+Early program-aided methods, such as PaL Gao et al.
 (2023) and PoT Chen et al.
-(2023) demonstrated significant gains by converting reasoning steps into executable programs, delegating precise computations to code interpreters.
+(2023), demonstrated significant gains by converting reasoning steps into executable programs, thereby delegating precise computations to code interpreters.
 Other lines of work, including WebGPT Nakano et al.
@@
 (2023) further generalized this approach by training language models to self-supervise API calls on various tasks such as arithmetic, translation, and retrieval.
-However, unlike DualDistill, these methods typically rely on short chain-of-thoughts primarily using prompting or heuristic-based tool invocation, lacking mechanisms for automatically balancing pure long reasoning against tool use based on task complexity.
+START Li et al.
+(2025b) and CoRT Li et al.
+(2025a) use hint-based prompting to activate tool use behavior, followed by rejection sampling fine-tuning (RFT) for self-improvement.
+However, unlike DualDistill, these methods typically focus on short-CoT, or primarily use prompting or heuristic-based tool invocation, lacking mechanisms to automatically balance long-term reasoning against tool use based on task complexity.
 Long Chain-of-Thought Reasoning.
-Recent approaches have highlighted significant performance improvements by scaling inference-time chain-of-thought length.
-OpenAI’s O1 Jaech et al.
+Recent approaches have highlighted significant performance improvements by scaling the length of the inference time chain-of-thought (CoT).
+GPT-o1 Jaech et al.
 (2024) and DeepSeek-R1 Guo et al.
-(2025) used outcome-driven reinforcement learning to generate extensive reasoning trajectories, substantially outperforming shorter-CoT baselines on complex math and reasoning benchmarks.
+(2025) used outcome-driven reinforcement learning to generate extensive reasoning trajectories, substantially outperforming short-CoT baselines in complex math and reasoning benchmarks.
 Similarly, S1 Muennighoff et al.
@@
 (2025).
-Nonetheless, long-CoT models frequently encounter overthinking, i.e, generating overly long reasoning that leads to redundant or incorrect outcomes, especially in tool use scenarios, in a phenomenon known as the reasoning-action dilemma Cuadron et al.
+However, long-CoT models frequently encounter overthinking, which is the generation of overly long reasoning that leads to redundant or incorrect outcomes, especially in tool use scenarios, a phenomenon known as the reasoning-action dilemma Cuadron et al.
 (2025).
@@
 (2025) and Search-R1 Jin et al.
-(2025) introduced reinforcement-learning-based retrieval policies within reasoning loops, achieving substantial performance improvements on open-domain question-answering tasks.
+(2025) introduced reinforcement-learning-based retrieval policies within reasoning loops, achieving substantial performance improvements in open-domain question-answering tasks.
 However, unlike these methods, DualDistill is specifically tailored for math tasks.
@@
 Shen et al.
-(2022) introduces self-distillation by using the outputs of the previous mini-batch as soft targets, improving output consistency without external teachers.
+(2022) introduces self-distillation using the output of the previous mini-batch as soft targets, improving the consistency of the output without external teachers.
 In the LLM era, fine-tuning LLMs on self-generated trajectories with a rule-based verifier has been applied in many domains, including coding tasks Jiang et al.
-(2024), formal proving Lin et al.
+(2024), theorem proving Lin et al.
 (2024), math tasks Kumar et al.